Yeast cells are to be separated from a fermentation broth. Assume that the cells are
spherical with diameter 5 µm and density 1.06 g/cm3. The viscosity of the culture
broth is 1.36 × 10-3 Ns/m2. At the temperature of separation, the density of the
suspending fluid is 0.997 g/cm3. 500 liters of broth must be treated every hour.
Specify Σ factor for a suitably-sized disk-stack centrifuge.
1
Expert's answer
2020-08-02T15:51:06-0400

Density of cells,p=1.06 g/cm3=1060 kg/m3 Diameter of cell, Dp=5m=5*10-6m Density of fluid,f=0.997 g/cm3=997 kg/m3 Rate of broth treated, Q=500 L/h=0.5 m3/h=1.39*10-4m3/s Viscosity